OpenStudy (anonymous):

a ferris wheel starts spinning at t=0s and stops at t=12s. If the ferris wheel made 5 loops during that time, what is the value for b in the equation?

OpenStudy (anonymous):

@wio can you help me?

OpenStudy (anonymous):

In what equation?

OpenStudy (anonymous):

i wasnt given one ......

OpenStudy (anonymous):

in radian measures please

OpenStudy (anonymous):

it is multiple choice, answer is either 2pi/5, 2pi, 2pi/12, or 5pi/6

OpenStudy (anonymous):

@mathstudent55 help me?

OpenStudy (anonymous):

A revolution is \(2\pi\). So 5 loops would be \(5\cdot (2\pi)=10\pi\)

OpenStudy (anonymous):

@Blooper9671111 You have left information out.